### v1

# Automated Bug Generation in the era of Large Language Models

Abstract.
Bugs are essential in software engineering; many research studies in the past decades have been proposed to detect, localize, and repair bugs in software systems. Effectiveness evaluation of such techniques requires complex bugs, i.e., those that are hard to detect through testing and hard to repair through debugging. From the classic software engineering point of view, a hard-to-repair bug differs from the correct code in multiple locations, making it hard to localize and repair. Hard-to-detect bugs, on the other hand, manifest themselves under specific test inputs and reachability conditions. These two objectives, i.e., generating hard-to-detect and hard-to-repair bugs, are mostly aligned; a bug generation technique can change multiple statements to be covered only under a specific set of inputs. However, these two objectives are conflicting for learning-based techniques: A bug should have a similar code representation to the correct code in the training data to challenge a bug prediction model to distinguish them. The hard-to-repair bug definition remains the same but with a caveat: the more a bug differs from the original code (at multiple locations), the more distant their representations are and easier to be detected.
We propose BugFarm, to transform arbitrary code into multiple complex bugs. BugFarm leverages Large Language Models (LLMs) to mutate code in multiple locations (hard-to-repair). To ensure that multiple modifications do not notably change the code representation, BugFarm analyzes the attention of the underlying model and instructs LLMs to only change the least attended locations (hard-to-detect). Our comprehensive evaluation of $320k+$ bugs from over $2.5M$ mutants generated by BugFarm and two alternative approaches demonstrates our superiority in generating bugs that are hard to detect by learning-based bug prediction approaches (up to 41% higher False Negative Rate and 11%, 6%, 29%, and 21% lower Accuracy, Precision, Recall, and F1 score) and hard to repair by state-of-the-art learning-based program repair technique (22% repair success rate compared to 34% and 49% of LEAM and $\mu$BERT bugs). BugFarm is efficient, i.e., it takes nine seconds for it to mutate a code without any prior training time.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXcopyright: noneprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Problem Statement and Significance. Machine learning is becoming the new automation technology, and software engineering automation is no exception. AI pair programming techniques such as Copilot (GitHub, 2023) are integrated into IDEs, and LLMs such as GPT-4 (OpenAI, 2023) and Bard (Google, 2023a) are now accessible through API. Following this trend, state-of-the-art software analysis techniques either fine-tune pre-trained code-language models or prompt LLMs to automate code-related tasks  [^1] We differentiate between LLMs and code-language models. The former models are large ($>10B$ parameters) and are used through zero- or few-shot prompting. The latter models are smaller and are used after fine-tuning.. This demands automated techniques for generating high-quality datasets to assess their true effectiveness. Concerning bug-related tasks such as bug prediction, localization, and repair, such datasets should include a diverse set of complex bugs, i.e., those that are hard to detect to challenge bug prediction techniques and hard to repair for debugging approaches.
From the classic software engineering perspective, hard-to-detect bugs often exist at locations reachable only under a particular combination of test inputs or edge cases (Lu et al., 2005; Böhme and Roychoudhury, 2014). A similar definition does not stand concerning learning-based bug prediction techniques. That is because they do not care what inputs trigger the bug in the given code. In contrast, they look for bug patterns in their training/fine-tuning data (Yang et al., 2015) or to check if the code representation is closer to buggy or correct examples they have seen (Ibrahimzada et al., 2022). As a result, to challenge the learning-based bug prediction techniques, one should inject unseen bug patterns such that the code representation of the generated bug and correct code are similar.
Hard-to-repair bugs usually involve multiple statements, making debugging techniques struggle to localize and understand their interactions (Li et al., 2018) to fix them automatically. A similar definition holds from the learning-based software engineering, but with a slight consideration: The power of machine learning is in learning from patterns that persist in their dataset. Consequently, modifying a code in multiple locations, where the learning-based debugging techniques specifically look to localize or repair the bug, would be counter-intuitive. Can existing techniques provide complex bugs, as defined, to evaluate learning-based techniques?
Research Gap. Several attempts have been made to construct real-world bug datasets, e.g., Defects4J (Just et al., 2014a), BugSwarm (Tomassi et al., 2019), BugsInPy (BugsInPy, 2023), RegMiner (Song et al., 2022), and ManySStuBs4J (sst, 2023). While representing real-world bugs, such datasets are only limited to the projects and bug patterns included in their mining efforts. More importantly, they may be included in the training/fine-tuning data of learning-based techniques, causing data leakage (Samala et al., 2020). To overcome these limitations, researchers have proposed automated bug generation techniques (Tian et al., 2022; Khanfir et al., 2023; Jia and Harman, 2009). Such techniques, by default, change a few locations in the code (easy to repair and detect) but can be configured to modify multiple lines. That said, changing multiple lines randomly (Khanfir et al., 2023) or following a set of heuristics (Tian et al., 2022) may result in mutants with a different code representation than the correct code, making them easy to detect. Furthermore, changing lines that learning-based repair techniques have been trained to look into and repair (Xia et al., 2023), such as conditional statements, results in bugs that can be easily repaired, even if they involve multiple lines.
Proposed Solution. To advance automated bug generation concerning the evaluation of learning-based bug-related tasks, we propose BugFarm. For a given arbitrary code, BugFarm prompts a Large Language Model (LLM) to mutate multiple statements. Given that LLMs are potentially creative in generative tasks, leveraging them helps avoid overfitting to a limited number of mutation operators and bug patterns. To ensure that changing too many locations does not drastically change the code representation of the mutant, BugFarm analyzes the attention of the underlying model and instructs LLMs to only change those the model attends least to. As a result, the generated code has a similar representation to the original one but differs within multiple locations. This makes them hard to detect by bug prediction approaches and hard to repair by debugging techniques. BugFarm is programming language-agnostic, i.e., making it a pragmatic approach to generate bugs for any programming language. Like related mutant generation techniques, BugFarm creates unconfirmed bugs, which should be validated through test execution. Our notable contributions are:
  • A new perspective into bug generation: We propose a novel technique for bug generation concerning the evaluation of the learning-based bug-related tasks. Our proposed technique, BugFarm, relies on the model’s attention to different parts of code to identify where to inject bugs so that the impact of changes on code representation is minimal. It leverages such information to craft contextual prompts for LLMs to mutate any arbitrary code. The implementation and artifact of BugFarm, including all the generated bugs, are publicly available (authors, 2023).
  • Empirical evaluation: We extensively evaluated BugFarm and two most recent learning-based bug generation approaches, LEAM (Tian et al., 2022) and $\mu$BERT (Khanfir et al., 2023). Our evaluation of $320k+$ bugs (from over $2.5M$ mutants) generated for $15$ Java projects answers research questions related to (1) the properties of generated bugs, (2) the ability of learning-based bug prediction models to detect them, (3) the effectiveness of a learning-based program repair technique to repair them, (4) the impact of prompt crafting, and (5) the performance characteristics of techniques. The results confirm that BugFarm bugs are hard-to-detect (up to 41% higher False Negative Rate and 11%, 6%, 29%, and 21% lower Accuracy, Precision, Recall, and F1 score) and hard-to-repair (22% repair success rate compared to 34% and 49% of LEAM and $\mu$BERT bugs). Also, generating effective bugs is more efficient using BugFarm.
2. Illustrative Example
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final E value : values) {
            dec.decorated().add(transformer.transform(value));
    return dec;
(a) Original Method
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if ((bag == null)) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final E value : values) {
            dec.decorated().add(transformer.transform(value));
    return dec;
(b) Bug generated by LEAM (Tian et al., 2022)
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final E value : values) {
            dec.decorated().add(transformer.transform(value));
    return dec;
(c) Bug generated by $\mu$BERT (Khanfir et al., 2023)
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        null;
    if (!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final E value : values) {
            dec.decorated().add(transformer.transform(value));
    return null;
(d) Bug generated by BugFarm
Figure 1. An illustrative example showing the bugs generated for the original code (a) using LEAM (b), $\mu$BERT (c), and BugFarm (d).
To illustrate the limitations of prior work and present the key idea behind BugFarm, we will use the code snippets in Figure 1. The original code in Figure 0(a) takes bag and transformer objects as inputs. If bag is not empty, it moves all the objects inside it to transformer. There are several ways to inject bugs into this code snippet by adding, deleting, or modifying the $13$ statements in its body (some statements are split into two lines for better presentation). These bugs are intended to challenge bug prediction and program repair models we used in our evaluation (section 5), which are fine-tuned on top of CodeT5 (Wang et al., 2021). Concerning the attention of the underlying model to the original code, the two least attended statements (details in algorithm 1) are highlighted in green.
We leverage three bug generation tools, namely, BugFarm, LEAM (Tian et al., 2022), and $\mu$BERT (Khanfir et al., 2023) to transform the original code snippets into bugs (bugs were confirmed through test execution). LEAM is a technique that represents code as a sequence of AST nodes and learns to apply grammar rules to select and modify the code for bug generation. $\mu$BERT selects code tokens corresponding to AST nodes, replaces them with a special token <mask>, and asks CodeBERT to replace them with new tokens for bug generation. BugFarm identifies the least attended statements and prompts an LLM to perform bug-inducing transformations only on those lines. Figures 0(b) and  0(c) show one of the generated bugs by LEAM and $\mu$BERT (LEAM and $\mu$BERT generate $19$ and $68$ mutants in total, out of which only $6$ and $1$ are confirmed bugs). BugFarm generates a bug shown in Figure 0(d). The lines changed to introduce a bug are highlighted in red.
BugFarm bug is notable from various perspectives. First, it involves multiple statements (hard to localize and repair). Second, both modified statements are among the least attended statements, making it hard for the model to distinguish the bug from the original code (hard to detect). LEAM and $\mu$BERT bugs modify only one line that is not among the least attended statements. As a result, when running against our studied bug prediction model (section 5.3) and program repair model (section 5.4), they were easily detected and repaired. In contrast, the same models failed to detect and repair BugFarm bug.
Figure 2. Overview of BugFarm
3. Approach Overview
Figure 2 provides an overview of BugFarm framework consisting of three major components: (1) Method Extractor, (2) Attention Analyzer, and (3) Bug Generator. BugFarm takes a project and a pre-trained code-language model as inputs and extracts the methods through a lightweight static analysis. For each method, it then identifies the bug location candidates by analyzing the model’s attention to individual code tokens. To generate bugs for each method, BugFarm crafts a prompt to an LLM, including the code to be transformed into bugs and additional contexts reflecting candidate locations for bug injection. From the generated bugs, BugFarm selects those that are syntactically correct, where the bug-inducing changes do not significantly change the attention of the model.
The Method Extractor component takes the input project and builds its corresponding parse tree to extract all the methods in the source files. These methods will be passed as an input to Attention Analyzer to identify the candidate locations for bug injection (more details in  section 4.1).
Recent state-of-the-art learning-based software engineering models are all transformer-based due to the naturalness of code (Hindle et al., 2016) and the promising results of such models in natural language processing (NLP) (Vaswani et al., 2017). Transformer-based models rely on attention for neural code representation. So, to identify the candidate locations to inject unnoticeable bugs into a given method, Attention Analyzer extracts the corresponding attention of the model to code tokens and identifies those with the lower contribution in the code representation, i.e., those with less attention weight values. Bug locations in BugFarm are at the statement level; thereby, BugFarm ranks the input method’s statements based on the #$\#T_{LA}/\#T$ values and chooses the ones with the lowest value. Here, $T_{LA}$ is the number of least attended tokens, and $T$ denotes the total number of tokens in the statement (more details in  section 4.2).
Finally, Bug Generator component takes the list of bug location candidates as input and creates a prompt consisting of natural language instruction, bug location candidates, and the method. It then sends the prompt to an LLM and collects the generated bugs in the response. For the generated bugs by LLM, BugFarm computes the extent to which the bug-inducing changes impact the model’s attention to the buggy method compared to the input method, and selects those with negligible impact on the attention. It also discards the bugs that are syntactically incorrect for the sake of quality (more details in section 4.3).
4. BugFarm
In this section, we will explain the details of three main components of BugFarm. Specifically, we will first discuss how BugFarm parses the subjects and extract methods as inputs for the Attention Analyzer component. Next, we will provide a background on the attention analysis required to understand the subsequent components. Finally, we answer two main questions, namely (1) How does BugFarm decide where to inject bugs?, and (2) how BugFarm generates a set of bugs?
4.1. Method Extractor
Our approach sets itself apart from most prior work, which relies on defect datasets for automated bug generation (Tian et al., 2022; Nong et al., 2023; Patra and Pradel, 2021). Instead, it focuses on injecting diverse and complex bugs into any well-maintained project using LLMs, regardless of the programming language. As a result, we assume that the input to the BugFarm is a project rather than a single method. As a first step, BugFarm first requires extracting the implemented methods in each input project. To that end, Method Extractor component leverages a language-specific parser depending on the programming language used in the input project, builds program parse trees, and extracts a list of methods and constructors. Next, it collects both signatures and body (excluding docstrings) using MethodDeclaration and ConstructorDeclaration blocks in the parse tree. Each method will be passed to the Attention Analyzer component, which we will describe next.
4.2. Attention Analyzer
State-of-the-art code-language models are based on the Transformer architecture (Vaswani et al., 2017)—an encoder to capture input’s contextual representations and a decoder to generate output tokens. To produce contextualized vector representation of a sequence of tokens, Transformers rely on Multi-Head Self-Attention. More specifically, for a method that consists of $n$ tokens $\overrightarrow{Tkn}=\{m_{0},\ldots,m_{n-1}\}$, a Transformer model with $\mathcal{L}$ layers takes $\overrightarrow{Tkn}$ as input and produces $\textbf{H}^{\ell}=[\textbf{h}_{0}^{\ell},...,\textbf{h}_{n-1}^{\ell}]$. Here, $\textbf{H}^{\ell}$ corresponds to the hidden vector representations in layer $\ell\in\{1,2,...,\mathcal{L}\}$. In each Transformer layer $\ell$, multiple self-attention heads are used to aggregate the outputs of the previous layer. Consequently, for each token $m_{i}$ in the method, the self-attention assigns a set of attention weights with respect to all other tokens in the input sequence, i.e., $Attention(m_{i})=\{\alpha_{i0},\dots,\alpha_{in-1}\}$, where $\alpha_{ij}$ indicates the relative attention of $m_{i}$ to $m_{j}$.
The hidden representation weights of Transformers will be computed based on the attention weights, making the attention weights crucial to identify the importance of each token in the final representation of the method. Recent research has shown that attention analysis can be utilized to interpret the knowledge learned during training (Hernández López et al., 2022; Wan et al., 2022; Ibrahimzada et al., 2022). Following their insights, BugFarm leverages self-attention weight analysis to scrutinize the input method and identify tokens (and subsequently, statements) with the lowest attention weights. These statements are where BugFarm can change during bug-inducing transformation without impacting the overall representation of the method. Algorithm 1 explains our approach for attention analysis, which takes the method to be transformed into a bug, threshold value $k$, and a transformer-based model as inputs and pinpoints the $k\%$ of the least attended statements $LAS$ as outputs. To that end, it first extracts the list of least attended tokens in the method $LAT$ (Lines 1-10) and uses them to pinpoint the least attended statements $LAS$ (Lines 11-20).
The algorithm first identifies the tokens $\overrightarrow{Tkn}$ and statements $\overrightarrow{Smt}$ in the given method and initializes the $LAT$ and $LAS$ variables to be empty (Lines 1-3). Next, it queries the model $M$ to extract the self-attention values (Line 4). For a model $M$ with $L$ layers and $H$ attention heads per layer, the attention values will be averaged across heads and layers, resulting in an $n\times n$ matrix, where $n$ is the number of tokens. For each token $m_{i}$ in the $method$, the algorithm further averages the attention weight relative to other tokens (averaging the values per column in the self-attention matrix) to compute a single attention weight value for each token $m_{i}$ in the method (Line 5). Given that we are interested in the least attended tokens in the code, Algorithm 1 sorts the attention weight vector, $\overrightarrow{TknAttnW}$, populates their corresponding indices in $\overrightarrow{SortedTknInd}$ (Line 6), and identifies the top $k\%$ of least attended tokens, $LAT$ (Lines 7-10).
With the least attended tokens extracted, the algorithm can identify the least attended statements, $LAS$. To that end, it weighs each statement by a score (Lines 11-14), which is the ratio of the number of least attended tokens in that statement, normalized by its length. The key idea here is that a statement with the highest amount of overlap between least attended tokens should achieve a lower score, thus, be considered the least attended statement. Without normalizing the statement length, longer statements will be penalized, and BugFarm never selects them as a bug injection location. Finally, the statements are sorted in ascending order based on their scores, and the least $k\%$ attended statements (we take the ceiling in case $k\%$ of total statements is less than one) will be returned as $LAS$ (Lines 15-20).
Inputs: Method $method$, Threshold $k$, Transformer-based model $M$
Output: Least attended statements $LAS$
[…]
7 $\mathit{i}\leftarrow 0$;
[…]
10       $\mathit{i}\leftarrow i+1$;
11
[…]
13 foreach $s_{i}\in Smt$ do
[…]
16
[…]
18 $\mathit{i}\leftarrow 0$;
[…]
21       $\mathit{i}\leftarrow i+1$;
22
return $LAS$
Algorithm 1 Attention Analyzer
4.3. Bug Generator
Bug generation involves modifying, adding, or deleting code segments from an original method to change the expected behavior of the program. The Bug Generator module of BugFarm takes a method and its corresponding set of $LAS$ identified in the previous step as inputs, and crafts LLM prompts to generate $N$ buggy versions for the method as outputs. Our intuition for making BugFarm configurable is that our bugs, very likely, will be used as training/fine-tuning data for bug-related tasks. So, generating multiple buggy versions of a single method would be helpful for a model to distinguish between buggy and non-buggy code easier. The total number of bugs, however, also depends on the size of the method and threshold value $k$. For example, $k=10\%$ for methods with less than $10$ statements returns one statement as LAS, and changing that statement in $N$ unique ways may be infeasible.
BugFarm’s prompts consist of three parts. The first part is the natural language instruction asking LLM to generate the bugs. The second part provides contextual information about where to inject the bug, i.e., only to consider the least attended statements, $LAS$, for making bug-inducing changes. Finally, we include the entire method, including both signature and the method body, in the prompt. Such prompts are LLM-agnostic, i.e., they can be used with various existing LLMs, such as LLaMa (Meta, 2023), PaLM (Google, 2023b), Copilot (GitHub, 2023), Alpaca (Taori et al., 2023), and ChatGPT (AI, 2023). The only consideration is to check if the size of the prompt matches the context window of the LLM.
Inputs: Method $method$, Least attended statements $LAS$, Number of bugs $N$, Transformer-based model $M$
Output: Buggy methods $Bugs$
[…]
6 foreach $Response\in Responses$ do
7       if $!isParseable(Response)$ then
8             $continue$;
9
[…]
11       $check\leftarrow true$;
12       foreach $stmt\in getDiff(Response,method)$ do
13             if $stmt\notin newLAS$ then
14                   $check\leftarrow false$;
15                   $break$;
16
17
18      if $check$ then
[…]
20
21
Algorithm 2 Bug Generator
After the LLM’s response to the prompts, Bug Generator component validates (1) if they are syntactically correct and (2) if the changes do not impact the attention of the model. The responses that do not pass these two checks will be discarded. Ideally, we want the produced bugs to be syntactically correct. To that end, our prompts also include instructions to produce syntactically correct code. However, our experiments showed that despite asking an LLM to produce syntactically correct code, specifically for longer methods, it generates responses that are not parsable. Furthermore, our goal is to generate bugs that have a very close neural representation with the original code. So, we assess the model’s attention to the generated bugs and only accept them if the changes are still among the least attended statements.
Algorithm 2 demonstrates BugFarm’s bug generation and validation approach. It starts by initializing the output variable $Bugs$ (Line 1) and getting the indices of LASs in $method$ (Line 2). Next, it adds an index to each statement in $method$ (Line 3). This will help us to refer to LASs in the prompt by number instead of including the whole statements, resulting in a reduction of the prompt size. This is specifically important for longer methods or statements, as the context window of LLMs is limited (Nguyen and Wong, 2023; Xu et al., 2022). Next, we will craft the prompt with the required context, i.e., indexed method and bug injection location, and send the prompt to an LLM (Lines 4-5)  [^2] The natural language part of our prompts is more complex than the example here. The readers can refer to our artifact to see the exact prompts (authors, 2023). Finally, once the LLM responds, we check if the generated bugs are syntactically correct (Lines 7-8) and if the changed statements are among the least attended statements by the model (Lines 9-14) to add them to the acceptable set of bugs (Lines 15-16).
5. Evaluation
To evaluate the effectiveness of BugFarm, we compare it with two state-of-the-art automated bug generation approaches, namely LEAM and $\mu$BERT, investigating the following research questions:
  1. RQ1:
    Characteristics of the Generated Bugs. Can BugFarm successfully inject bugs into any arbitrary code? What are the characteristics of the generated bugs by each approach? To what extent do our bugs overlap with those generated by other approaches?
  2. RQ2:
    Effectiveness in Generating Hard-to-Detect Bugs. How well do learning-based bug prediction models perform at detecting BugFarm bugs compared to other techniques?
  3. RQ3:
    Effectiveness in Generating Hard-to-Repair Bugs. To what extent can a learning-based Automated Program Repair (APR) technique repairs BugFarm bugs compared to that generated by alternative approaches?
  4. RQ4:
    Necessity of Prompt Engineering. What are the benefits of prompt engineering in BugFarm?
  5. RQ5:
    Performance. How long does it take to generate and validate bugs using BugFarm and other approaches?
5.1. Experimental Setup and Data Availability
Alternative Approaches. We compare BugFarm with two most recent mutant generation techniques, $\mu$BERT (Khanfir et al., 2023) and LEAM (Tian et al., 2022). To inject mutants, $\mu$BERT selects AST nodes representative of program behavior—literals, identifiers, expressions, assignments, object fields, method calls, array access, and types. Then, it replaces the tokens in selected AST nodes with the spacial token <mask> and uses CodeBERT to predict the masked token. The intuition is that if CodeBERT predicts a token different from the original one, the transformation introduces a bug. LEAM is a deep learning-based technique that learns to mutate code from large examples of real-world bugs. To that end, they represent code as a sequence of AST nodes and learn to apply eight grammar rules to select and modify the code. Both $\mu$BERT and LEAM claim to generate better bugs (mutants confirmed by test execution) compared to classic approaches, namely PIT (Coles et al., 2016) and Major (Just, 2014). Therefore, we did not include classic techniques in our evaluation.
For BugFarm, we used the threshold value $k=10\%$, mainly because alternative approaches do not change more than a handful of statements in a given code (more details in section 5.2). Our experimental results in the rest of this section confirm that even with such a low threshold, we still surpass other techniques. We expect a higher threshold will make our bugs more complex compared to other approaches, improving margins. We also configured BugFarm to generate at most three mutants per method ($N=3$ in Algorithm 2) due to budget allowance. The current implementation of BugFarm’s Bug Generator component uses GPT-3.5-turbo as an LLM due to its superiority over a large number of models (Bubeck et al., 2023; Zhang et al., 2023; OpenAI, 2023; Pan et al., 2023) and cost-effectiveness. Using a more advanced LLM such as GPT-4 likely yields better results, not worse.
Bug Validation. Alternative approaches we compare against generate mutant, which may not necessarily change the outcome of test execution. Worse, the mutants may not even be compilable. For BugFarm, although we ask the LLM to generate syntactically correct bugs, it is still not guaranteed that the generated code executes and changes the outcome of test execution. This entails implementing a validation process to confirm if the generated mutants/bugs are compilable and can change the outcome of test execution to consider them bugs. Our validation procedure follows these steps: We first run existing test suites on the original code and select the green tests. Next, we take all the generated mutants/bugs by different techniques, compile them, and run the previously green tests against them (each mutant will be tested in isolation). We discard those that do not compile or result in no failure after test execution. All the experiments answering RQ1-RQ5 are done with confirmed bugs.
Bug Prediction Models. When selecting bug prediction models, we had to consider the following criteria: (1) the current implementation of BugFarm’s Attention Analyzer component works on transformer-based models, including customized trained models and those fine-tuned on pre-trained code-language models; (2) we require the availability of models to perform the attention analysis, so closed-source models such as Codex (Chen et al., 2021) are off the table; (3) to compare against other bug prediction approaches, we require models with either pre-trained models or required artifacts for training available. We could not find any custom bug prediction model publicly or per request available with such characteristics. Consequently, we chose three pre-trained code-language models that are widely used by the research community, namely CodeBERT (Feng et al., 2020), CodeT5 (Wang et al., 2021), and NatGen (Chakraborty et al., 2022) [^3] Fine-tuning larger models requires non-trivial computing resources. Our experiments will show that models superior in other code-related tasks (e.g., CodeT5 over CodeBERT) show the significance of BugFarm better. We expect this to hold for larger models as well.. We fine-tuned these models using the real world and synthetic bugs, i.e., those generated by LEAM and $\mu$BERT, and compared their effectiveness in predicting our bugs.
CodeBERT is an encoder-only transformer model based on BERT (Devlin et al., 2018) architecture, which is trained on $2.1M$ bi-modal (natural language and code pairs) and $6.4M$ uni-modal (code only) data from CodeSearchNet (Husain et al., 2019) dataset. The main learning objectives in CodeBERT are Masked Language Modeling (MLM)—the model learns to predict the tokens replaced by a special mask token—and Replaced Token Detection (RTD)—the model learns to detect which token does not belong to the original data. CodeT5 is an encoder-decoder transformer model based on T5 (Raffel et al., 2020) architecture. CodeT5 is trained on $8.35M$ functions from various programming languages provided by CodeSearchNet (Husain et al., 2019) and BigQuery (Big, 2023) dataset, and its training objectives include masked span and masked identifier prediction. Such objectives enable the model to understand code semantics better than CodeBERT.
NatGen is also an encoder-decoder transformer model, trained on a generative task of naturalizing source code. Specifically, NatGen starts with CodeT5—based model—parameters and continues the training with a new objective, i.e., re-constructing the original code (natural) given transformed code (de-natural). It uses $8.1M$ pairs of natural and de-natural functions from CodeSearchNet (Husain et al., 2019) and C/C#. We used the base model of CodeBERT, CodeT5, and NatGen for our experiments.
Automated Program Repair Model. To evaluate the effectiveness of learning-based techniques in repairing generated bugs, we used FitRepair (Xia et al., 2023). FitRepair is a state-of-the-art APR technique that outperforms existing approaches. It leverages information retrieval and static analysis to implement domain-specific fine-tuning and prompting strategies. We used FitRepair with CodeT5-Large in a zero-shot manner to generate patches for $\mu$BERT, LEAM, and BugFarm bugs and validated if the patches thoroughly fix the bugs through test execution.
Subjects. BugFarm is programming-language agnostic; none of its components depend on a specific programming language. However, we chose Java projects in our experiments for the following reasons: (1) LEAM’s pre-trained model is on Java, and they have no alternative training dataset for other programming languages; (2) we needed real-world bug datasets for RQ2, and most of the existing real-world bug datasets are in Java. For a fair comparison, we used Defects4J $V2.0$ projects as a baseline for bug generation since the alternative approaches are shown to work on them with no issues. The current version of BugFarm supports Maven projects only, so we excluded Mockito and Closure projects from the subjects. The first two columns of Table 1 show the list of our $15$ subjects and the number of methods per subject used for bug generation.
Table 1. Comparing the characteristics of bugs generated by BugFarm, $\mu$BERT, and LEAM. M: # Methods, UB: # Unconfirmed Bugs, CB: # Confirmed Bugs, SI: # Statements Involved in bug generation, LD: Lines Deleted in bug generation, ED: Edit Distance, OL: Overlap with LEAM, OM: Overlap with $\mu$BERT.
Subjects M BugFarm LEAM (Tian et al., 2022) $\mu$BERT (Khanfir et al., 2023)
UB CB SI LD ED OL OM UB CB SI LD ED UB CB SI LD ED
cli 276 438 263 (60%) 2.86 0% 32.14 ¡14%,0.64¿ ¡8%,0.67¿ 3959 1375 (34.73%) 2.56 13.82% 26.2 12641 1528 (12%) 2.00 0.07% 19.38
codec 901 1586 357 (23%) 3.12 0% 32.39 ¡3%,0.67¿ ¡4%,0.85¿ 13603 2114 (16%) 2.59 7.95% 20.33 42529 9541 (22%) 2.01 0% 17.89
collections 4440 7373 2958 (40%) 2.55 0% 27.01 ¡7%,0.67¿ ¡5%,0.73¿ 56196 9356 (17%) 2.71 8.87% 21.33 132009 21367 (16%) 2.01 0% 20.02
compress 4123 6544 464 (7%) 3.04 0% 30.29 ¡5%,0.69¿ ¡2%,0.80¿ 61488 9085 (15%) 2.57 12.87% 24.56 167259 22657 (14%) 1.97 0% 23.83
csv 248 379 250 (66%) 2.66 0% 24.30 ¡12%,0.66¿ ¡1%,0.75¿ 3560 1181 (33%) 2.83 14.65% 33.60 818 106 (13%) 1.98 0% 6.81
jxpath 1672 2601 912 (35%) 3.11 0% 26.09 ¡6%,0.69¿ ¡5%,0.78¿ 26615 5645 (21%) 2.86 16.44% 26.81 90153 10678 (12%) 2.00 0% 25.88
lang 3810 6505 3734 (57%) 2.95 0% 29.76 ¡7%,0.68¿ ¡3%,0.76¿ 54101 15861 (29%) 2.66 10.01% 21.60 80200 15618 (19%) 2.00 0.02% 21.13
math 5796 8935 6104 (68%) 3.27 0% 33.12 ¡4%,0.67¿ ¡2%,0.77¿ 92243 40299 (44%) 2.70 8.56% 17.15 359023 83206 (23%) 2.01 0.01% 18.30
gson 985 1497 537 (36%) 3.19 0% 36.43 ¡8%,0.7¿ ¡2%,0.68¿ 14496 2957 (20%) 2.60 18.16% 27.89 25359 2067 (8%) 2.00 0.05% 19.84
jackson-core 2626 4494 1707 (38%) 3.85 0% 41.85 ¡3%,0.68¿ ¡2%,0.72¿ 44364 9697 (22%) 2.43 22.74% 25.74 62963 9296 (15%) 2.09 0% 21.58
jackson-db 8076 12812 3855 (30%) 3.76 0% 41.70 ¡6%,0.66¿ ¡5%,0.74¿ 115859 17789 (15%) 2.78 12.61% 28.96 254947 22300 (9%) 2.01 0% 22.38
jackson-xml 586 877 249 (28%) 3.80 0% 44.24 ¡6%,0.68¿ ¡5%,0.71¿ 9935 1541 (16%) 2.67 13.17% 31.46 16420 1225 (7%) 1.98 0.08% 14.64
jfreechart 8602 13679 3035 (22%) 2.88 0% 25.26 ¡5%,0.71¿ ¡0.6%,0.81¿ 164460 16716 (10%) 2.44 11.23% 19.64 111341 7026 (6%) 2.01 0.07% 28.20
joda-time 4279 7833 3574 (46%) 2.38 0% 20.61 ¡6%,0.66¿ ¡2%,0.74¿ 69548 13739 (20%) 2.87 7.01% 24.73 141682 28010 (20%) 2.00 0% 21.21
jsoup 1642 2694 1335 (50%) 2.79 0% 28.47 ¡9%,0.66¿ ¡4.5%,0.7¿ 22632 5311 (23%) 2.52 14.54% 23.54 30420 4002 (13%) 2.24 0.05% 21.07
Total 48062 78247 29334 (37%) - - - - - 753059 152666 (20%) - - - 1527764 238627 (16%) - - -
Average 3204 5216 1956 (-) 3.08 0% 31.58 ¡7%,0.67¿ ¡3.49%,0.75¿ 50204 10178 (-) 2.65 13% 24.90 101851 15908 (-) 2.00 0.02% 20.14
Evaluation Metrics. To compare the performance of bug prediction models, we use accuracy, precision, recall, F1 score, False Positive Rate (FPR), and False Negative Rate (FNR) as our metrics. To evaluate the APR results, we measure the repair success rate, i.e., the number of bugs the technique successfully patches. We define TP, TN, FP, and FN as below:
  • True Positive (TP). The code is buggy; the model predicts it as buggy.
  • True Negative (TN). The code is not buggy; the model predicts it as non-buggy.
  • False Positive (FP). The code is not buggy; the model predicts it as buggy.
  • False Negative (FN). The code is buggy; the model predicts it as non-buggy.
Data Availability. The implementation of BugFarm and all the artifacts required for reproducing the results presented in this paper are publicly available (authors, 2023).
5.2. RQ1: Quality of the Generated Bugs
To answer this research question, we compared the characteristics of the bugs generated by BugFarm with alternative approaches using the metrics listed below. The metrics are computed for all the methods and averaged for each project. For BugFarm, we have also averaged the numbers across all the models (CodeBERT, CodeT5, and NatGen) [^4] BugFarm data per base model is publicly available in the artifact website..
  • # Unconfirmed bugs: This metric, computed under columns UB in Table 1, shows the total number of unconfirmed bugs/mutants generated for each project. All the approaches successfully generate mutants/bugs for all the methods in the subject projects. $\mu$BERT generates $1.5$ and $8$ times more bugs compared to LEAM and BugFarm, respectively.
  • # Confirmed bugs: This metric, shown under columns CB in Table 1, measures the number of confirmed bugs per each project generated by different approaches. The numbers inside the parenthesis indicate the percentages of confirmed bugs compared to unconfirmed bugs. A higher percentage implicitly implies the superiority of the technique in bug generation, i.e., the code transformations are more likely to be bugs. From these results, we see that 37% of the generated bugs by BugFarm are real bugs, compared to 20% and 16% for LEAM and $\mu$BERT, respectively. Except for one project (Commons Compress), the #CB/#UB ratio for BugFarm is higher, with a significant margin. The rest of the metrics are computed for confirmed bugs, which we refer to as bugs for simplicity.
  • # Statements involved in bug generation: This metric indicates the number of statements added, removed, or modified to generate the bugs. On average, even with the threshold values of $k=10\%$, BugFarm changes more statements to generate bugs compared to LEAM and $\mu$BERT. Although #SI is higher for BugFarm bugs, given that these statements are among the least attended statements, we will later see that the models will have a harder time distinguishing them from the correct code (more details in section 5.3).
  • # Lines deleted: Deleting an entire statement, while likely to create a bug, may not be the best bug-inducing transformation, as it will change attention significantly and result in a runtime exception rather than a semantic bug. So, we also wanted to compare the generated bugs from this aspect. Columns LD in Table 1 show the percentage of bugs created with only lines deleted (or commented) during bug generation. BugFarm does not delete (or comment) any statement through bug-inducing transformation ($0\%$ on average for all the projects), compared to LEAM ($13\%$) and $\mu$BERT ($0.02$%).
  • Edit distance: We also wanted to compare the edit distance between the original and buggy versions generated by each technique. We used Levenshtein (Levenshtein et al., 1966) edit distance, which measures the minimum number of single-character edits—insertions, deletions, or substitutions—required to change one string into another. Each code is represented as a string of characters to compute the Levenshtein edit distance. The results for edit distance normalized by #CB are available under columns ED in Table 1. Compared to LEAM and $\mu$BERT, BugFarm’s bugs have higher ED values. This again indicates that BugFarm’s changes to the code are bigger yet less noticeable, as we will show in RQ2 and RQ3. Between $\mu$BERT and LEAM, the latter generates bugs different from the original code in more places.
  • Uniqueness: Finally, we were interested to see how much our bugs overlap with those generated by $\mu$BERT and LEAM. To that end, we measured the Exact Match (EM) and CodeBLEU (Ren et al., 2020) values between each BugFarm bug with all the corresponding bugs generated by $\mu$BERT and LEAM. For example, for a given method A, if BugFarm generates three bugs ${b_{1},b_{2},b_{3}}$ and LEAM generates four ${l_{1},l_{2},l_{3},l_{4}}$, we construct $12$ pairs of $\langle b_{i},l_{j}\rangle$ to compute the EM and CodeBLEU. EM is a strict all-or-nothing metric; being off by a single character results in a score of 00. If the characters of $b_{i}$ exactly match the characters of $b_{j}$, EM = $1$ for the pair; otherwise, EM = 00. CodeBLEU is a metric to measure weighted n-gram match between the pairs by considering not just the code tokens but also code syntax via abstract syntax trees (AST) and code semantics via data flow.
    The EM and CodeBLUE values are shown under columns OL (overlap with LEAM) and OM (overlap with $\mu$BERT) in Table 1. The first number indicates EM, and the second one is CodeBLEU. These numbers show that only $7\%$ and $3.49\%$ of the total bugs generated by BugFarm for all the projects overlap with LEAM and $\mu$BERT, respectively, confirming the uniqueness of BugFarm bugs. The average CodeBLEU values are also $0.67$ and $0.75$. The high number for CodeBLEU is not a threat here, considering bug transformations do not greatly change the code compared to the original one. In fact, it shows that although BugFarm bugs are better at challenging learning-based techniques, they are semantically similar to LEAM and $\mu$BERT bugs; potentially as effective as them in their evaluated tasks (Tian et al., 2022).
{tcolorbox}
[enhanced, frame hidden, boxsep=0pt]Summary. Compared to other techniques, BugFarm’s bug-inducing transformations involve more statement modification and change of several code tokens. The overlap between BugFarm bugs and other approaches is low, demonstrating their uniqueness.
5.3. RQ2: Effectiveness in Generating Hard-to-Detect Bugs
In this research question, we investigate the effectiveness of bug prediction models on BugFarm bugs compared to alternative approaches. Given the unavailability of off-the-shelf models as discussed in section 5.1, we fine-tuned three pre-trained code-language models (CodeBERT, CodeT5, and NatGen) using real-world and synthetic bugs. We adapted best practices for fine-tuning transformer models and used the same class distribution in fine-tuning, validating, and testing bug prediction models. All models were fine-tuned for at most $10$ epochs with loss-based early-stopping criteria of two consecutive epochs 555this is the default setting for fine-tuning CodeT5. To ensure a fair comparison, we checked that all the models converged before $10$ epochs. and selected the model with the least validation error in our experiments. For more reliable results, we repeated the evaluation 10 times and reported the average values.
Generally, one should run BugFarm on each fine-tuned model to generate bugs accordingly. However, our investigation showed that fine-tuning does not greatly impact the set of LAS (Algorithm 1 in section 4.2). This is consistent with the findings of prior work that shows fine-tuning only changes the attention of the last few layers, not impacting the overall attention of the model (Shi et al., 2023). It also shows that many of the generated BugFarm bugs in this study are reusable by other researchers. Consequently, we only generated bugs for the methods per each model whose LAS set was changed compared to the baseline pre-trained model.
5.3.1. Fine-tuning on synthetic bugs.
We fine-tuned the three baseline pre-trained models on LEAM and $\mu$BERT bugs. This provided us with six fine-tuned bug prediction models, namely, $\mu$BERT-CodeBERT, $\mu$BERT-CodeT5, $\mu$BERT-NatGen, LEAM-CodeBERT, LEAM-CodeT5, and LEAM-NatGen. Our goal is not only to evaluate these models in predicting BugFarm bugs, but also to see the extent to which BugFarm bugs are harder to be detected than other techniques. To that end, we evaluated the fine-tuned bug prediction models on BugFarm and the other approach, whose bugs were not used for fine-tuning. For example, we evaluated $\mu$BERT-CodeBERT on bugs generated by BugFarm and LEAM. To ensure a fair comparison, we considered methods per each project that both comparing techniques had generated confirmed bugs for them. This will result in the same FPR values for both techniques, but allows us to focus on the effectiveness of bugs for evaluating the model better. This is important from the security perspective since missing a bug/vulnerability is more severe than marking a correct code as buggy (Ami et al., 2023).
The rows 1–12 in Table 2 show the result of this experiment. For Accuracy, Precision, Recall, and F1 score, the lower metric value indicates the approach’s superiority (bugs are harder to distinguish from correct code, hence being detected). For FNR, a higher metric value indicates the bugs are harder to detect. From these results, we can clearly see that BugFarm bugs always achieve higher values for FNR (average margin of 23% (min=2.76%,max=40.89%)) and lower values for the other metrics (average margin of Accuracy=5.97%, Precision=1.48%, Recall=15.5%, and F1-score=10.78%). Furthermore, by looking at the F1 score values, we observe that the models fine-tuned on LEAM have a harder time detecting BugFarm bugs than those fine-tuned on $\mu$BERT. We believe this is because $\mu$BERT bugs are more diverse due to changing many tokens of the code and combining them through beam search, compared to LEAM bugs that try to mimic the bugs in their dataset, scrapped from the GitHub issue trackers.
5.3.2. Fine-tuning on real-world bugs.
To avoid any bias in our conclusion based on synthetic bugs, we also fine-tuned baseline pre-trained models with real-world bugs from three datasets, namely Defects4J (Just et al., 2014a) [^6] We could use the bugs from Mockito and Closure projects, which were excluded from our subjects since BugFarm only supports Maven at this time, BugSwarm (Tomassi et al., 2019), and RegMiner (Song et al., 2022). The real-world evaluation dataset consists of $723$, $3285$, and $36412$ original and buggy methods from Defects4J, BugSwarm, and RegMiner, respectively. This resulted in three bug prediction models, i.e., REAL-CodeBERT, REAL-CodeT5, and REAL-NatGen. We evaluated each model on BugFarm, $\mu$BERT, and LEAM bugs. The rows 13–21 in Table 2 show the results of this experiment, with margins indicating the difference with respect to second-best synthetic bug dataset. These results show similar trends we observed with models fine-tuned on synthetic bugs, i.e., BugFarm bugs result in higher FNR (54.35% on average) and lower Accuracy (50.86% on average), Precision (51.52% on average), Recall (45.64% on average), and F1 score (47.34% on average) values.
We can also see that models fine-tuned on real-world bugs underperform those fine-tuned on synthetic bugs across all metrics. A possible justification for this observation is the distribution shift between real-world bugs (used for fine-tuning) and synthetic bugs (used for testing). The two root causes for these distribution shifts are (1) the real-world bugs belonging to projects different than those from which we generated the synthetic bugs; and (2) the nature of real-world bugs is different from synthetic bugs. Especially for BugFarm and $\mu$BERT, the bug generation objective does not include any similarity to real-world bugs.
Table 2. Effectiveness of studied fine-tuned models in predicting synthetic bugs. Each distinct set of rows show the same pre-trained model that is fine-tuned on the same dataset, but tested on different bug datasets.
Fine-tune-Model-Test Acc Prec Rec F1 FPR FNR
1 $\mu$BERT-CodeBERT-BugFarm 71.95 (4.40% normal-$\downarrow$) 90.04 (1.17% normal-$\downarrow$) 49.35 (11.83% normal-$\downarrow$) 63.76 (8.05% normal-$\downarrow$) 5.46 (0.00% --) 50.65 (15.04% normal-$\uparrow$)
2 $\mu$BERT-CodeBERT-LEAM 75.26 91.11 55.97 69.34 5.46 44.03
[…]
4 LEAM-CodeBERT-$\mu$BERT 77.92 95.75 58.43 72.57 2.59 41.57
[…]
6 $\mu$BERT-CodeT5-LEAM 76.08 87.9 60.48 71.66 8.33 39.52
[…]
8 LEAM-CodeT5-$\mu$BERT 78.4 91.43 62.66 74.36 5.87 37.34
[…]
10 $\mu$BERT-NatGen-LEAM 75.43 85.44 61.3 71.38 10.44 38.7
[…]
12 LEAM-NatGen-$\mu$BERT 76.72 90.91 59.39 71.84 5.94 40.61
[…]
14 REAL-CodeBERT-$\mu$BERT 55.22 58.4 36.31 44.78 25.87 63.69
15 REAL-CodeBERT-LEAM 57.68 61.44 41.23 49.35 25.87 58.77
[…]
17 REAL-CodeT5-$\mu$BERT 51.5 51.57 49.42 50.47 46.42 50.58
18 REAL-CodeT5-LEAM 53.79 53.77 53.99 53.88 46.42 46.01
[…]
20 REAL-NatGen-$\mu$BERT 50.85 50.71 61.16 55.45 59.45 38.84
21 REAL-NatGen-LEAM 53.82 53.02 67.1 59.23 59.45 32.9
{tcolorbox}
[enhanced, frame hidden, boxsep=0pt]Summary. Bug prediction models, regardless of whether the fine-tuning dataset is from a synthetic or real-world bug dataset, have a harder time detecting BugFarm bugs than other synthetic bugs.
5.4. RQ3: Effectiveness in Generating Hard-to-Repair Bugs
Table 3. Effectiveness of FitRepair in repairing synthetic bugs. SI: Statements involved.
LEAM (Tian et al., 2022) $\mu$BERT (Khanfir et al., 2023) BugFarm-CodeBERT BugFarm-CodeT5 BugFarm-NatGen
Total Bugs (SI=1,SI=2,SI¿2) 100 (84,15,1) 100 (92,8,0) 100 (52,26,22) 100 (34,25,41) 100 (50,16,34)
[…]
To further demonstrate the complexity of our bugs, we evaluate the ability of FitRepair in repairing BugFarm, $\mu$BERT, and LEAM bugs. The default configuration of FitRepair aims to generate $5000$ patches for a single bug and will terminate if this takes more than five hours. Given the size of our subject bugs and the non-negligible patch validation time, we reduced the maximum number of patches to 100100100100777This is the max allowance, and the tool often generates a lower number of patches (confirmed by the developers of FitRepair that this does not result in drastic performance degradation) and evaluated it on $100$ bugs sampled from each bug dataset. When sampling, we controlled for selecting bugs from the same methods for all approaches. To select the methods, we sorted them based on the descending order of method size—measured by the number of characters—and picked the top $100$. Our rationale is that the potential locations for bug injection increase when methods are longer. As a result, the likelihood of observing different bugs produced by each technique is higher (similar to the illustrative example in section 2). The collected bugs are from $11$ out of $15$ subject projects.
The first row in Table 3 shows the distribution of generated bugs based on the number of statements involved in bug generation (#SI from RQ1). Most subject bugs from LEAM and $\mu$BERT differ in only one line with the original code, while BugFarm bugs are more diverse concerning this metric. Each bug dataset took FitRepair two to seven hours to generate $35$ patches, on average. We validated all the generated patches for each bug, and if one of the patches results in a green test suite, we ignore other patches and consider the bug repaired by FitRepair. The validation process of $500$ bugs took over $50$ hours. The second row of Table 3 shows the percentages of bugs that FitRepair successfully patched.
FitRepair successfully repaired 34% and 49% of the LEAM and $\mu$BERT bugs, respectively. In contrast, it can only repair 27%, 22%, and 23% of BugFarm bugs generated for each pre-trained baseline. This is not surprising, as APR techniques are known to perform better in repairing bugs with SI=1. In fact, 88% of correct patches for $\mu$BERT and LEAM only differ with the corresponding bug in one line. This value is only 64% for BugFarm, which implies a higher complexity of BugFarm’s one-line bugs compared to alternative approaches. Our deeper investigation of the nature of bugs shows that LEAM and $\mu$BERT bugs mostly change the conditional/branch statements. Since repair templates of FitRepair are specifically designed to repair bugs in conditional/branch statements (logical expression in if statement), it can repair LEAM and $\mu$BERT bugs better. In contrast, BugFarm is more creative in introducing bug injection due to the power of LLMs in code synthesis, and it considers locations that the model attends less to for injecting the bugs. Consequently, it can challenge learning-based APR techniques crafted to repair known bug patterns. By looking at the other patches that fixed bugs with SI¿=2, we observed the same pattern, i.e., there was at least one statement changing the conditional statements.
{tcolorbox}
[enhanced, frame hidden, boxsep=0pt]Summary. When applied to the same method, BugFarm generates bugs that are harder to repair by learning-based repair techniques compared to alternative approaches. The power of BugFarm will become more evident when the methods are longer, letting it change multiple locations in the method to introduce the bug.
5.5. RQ4: Necessity of Prompt Engineering
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final E value : values) {
            dec.decoratedBag().add(transformer.transform(value));
    return dec;
(a) Incorrect method call
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final Object value : values) {
            dec.decorated().add(transformer.transform(value));
    return dec;
(b) Incorrect loop variable type
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!bag.isEmpty()) {
        // Removed the array initialization
        bag.clear();
        for (final E value : values) {
            dec.decorated().add(transformer.transform(value));
    return dec;
(c) Removed array initialization
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final E value : values) {
            // Removed the method call within the loop
    return dec;
(d) Removed method call
public static <E> List<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final E value : values) {
            dec.decorated().add(transformer.transform(value));
    return dec;
(e) Incorrect return type
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final Object value : values) {
            dec.decoratedBag().add(transformer.transform(value));
    return dec;
(f) Combination of bug 2(a) and 2(b)
Figure 3. The bugs generated without prompt engineering for the code in Figure 0(a).
The quality of prompts significantly impacts the quality of the model’s response (Liu et al., 2023); the more context you provide to the model in the prompt, the better the response it produces. In the context of our problem, identifying the location of Least Attended Statements (LASs) brings three benefits: (1) generating bugs with a similar representation compared to the original code (hard to detect bug), (2) enforcing the model to change multiple lines (hard to repair bug), and (3) reducing the search space, i.e., the location in which the bug can be injected. This will help the model focus on the most important part of the code and generate fewer yet more effective bugs.
Our previous research questions confirm the ability of prompt engineering to generate complex bugs. Generating those bugs without prompt engineering will require more time and cost more. For example, consider prompting an LLM to generate a bug for the code illustrated in Figure 0(a), implementing a logic through $13$ lines of code (some lines are split for better presentation). There are two least attended statements (LASs) in this method, highlighted in green. The probability that LLM changes one of them is $\frac{2}{13}=0.15$, i.e., with a probability of $15\%$, LLM can generate a hard-to-detect bug. However, to ensure the bug is also hard to repair, LLM should ideally modify the two highlighted lines. The probability of LLM choosing the two statements highlighted in green (the LASs) is only 1%, $\frac{1}{\binom{13}{2}}=0.01$. The code snippets in Figure 3 show different bugs generated by prompting GPT-3.5-turbo six times (buggy lines highlighted in red). We used a simple, not crafted prompt, asking it to generate a compilable bug from the given method. We can see that the LLM is quite creative in injecting different bugs (statement deletion, changing types, and incorrect method calls) into the code. But none of the responses change any of the least attended statements and, hence, have no overlap with the BugFarm bugs.
In general, for a given method with $n$ statements, the bug-inducing change can be applied to any line. BugFarm selects only $k$ percent of statements to be changed and instructs LLM to modify them, reducing the search space greatly. With the threshold value $k=20\%$, we can reduce the search space by $80\%$. Without attention analysis, a naive approach for search space reduction would randomly select $k$ percent of the statements. If not among the least attended statements, they can be easily detected or repaired by learning-based techniques, as shown in previous research questions.
{tcolorbox}
[enhanced, frame hidden, boxsep=0pt]Summary. Prompt crafting help BugFarm in generating hard-to-detect and hard-to-repair bugs. The chance of generating a similar dataset without prompt engineering is very narrow, if not impossible, due to the relatively large search space for moderate to long methods.
Figure 4. Performance of BugFarm compared to alternative approaches in unconfirmed bug generation
5.6. RQ5: Performance
To assess the performance characteristics of BugFarm, we measured the time required to extract attention weights from the models and the time it takes to prompt the LLM (GPT-3.5-turbo). We also compared the total time for bug generation in BugFarm compared to alternative approaches. All the experiments were performed on a workstation with NVIDIA GeForce RTX $3090$ GPUs ($24$GB GDDR6X memory) and $24$ 3.50GHz Intel 10920X CPUs ($128$ GB of memory). Figure 4 shows the results, where the red dashed line indicates the mean value. Looking at Attention Analyzer in Figure 4, we can see that it takes $68$, $69$, and $86$ milliseconds on average for BugFarm to extract and analyze attention weights from CodeBERT, CodeT5, and NatGen models. Prompting the LLM for every method also takes about $9$ seconds on average (Prompting box chart at the middle). The prompting time can be affected by multiple factors, including the traffic on the model and prompt size, i.e., the number of tokens in the prompt.
We also measured the total time each technique takes to generate unconfirmed bugs per method (End-to-End Unconfirmed Bug Generation in Figure 4). Compared to attention analysis and prompting, the overhead of other sub-components in BugFarm, such as parsing and bug selection, is negligible. Therefore, the average time for generating all the BugFarm unconfirmed bugs per each method is $9.29$ seconds (mostly dominated by the prompting time). In comparison, it takes LEAM and $\mu$BERT $35$ and $28$ seconds on average to generate unconfirmed bugs per method. This is because these approaches generate more unconfirmed bugs, as illustrated in Table 1 under #UB columns. For LEAM, this time does not include the training time of the model, which is $24$ hours  [^8] This number is quoted from the paper.. Also, since $\mu$BERT takes a long time to mutate big classes, we put a timeout of $15$ minutes on it to avoid a long generation time. Validation of unconfirmed bugs was also very time-consuming; we spent around $80,000$ CPU core-hours to validate over $2.5$ million mutants from $\mu$BERT, LEAM, and BugFarm.
{tcolorbox}
[enhanced, frame hidden, boxsep=0pt]Summary. Compared to alternative approaches, BugFarm is an efficient and scalable technique for generating bugs. More precisely, BugFarm is 74% and 67% faster in end-to-end unconfirmed bug generation than LEAM and $\mu$BERT, respectively.
5.7. Discussion
Our comprehensive evaluation of BugFarm compared to the two most recent alternative approaches confirms the ability of generated bugs to challenge learning-based bug prediction and repair techniques. We want to emphasize that our objective in this paper is to generate hard to detect and repair bugs, concerning learning-based techniques. As a result, comparing with real-world bugs is out of the scope of this paper. We have no claim that BugFarm bugs mimic real-world bugs (because they do not need to) or that they do not. In fact, BugFarm leverages GPT-3.5-turbo for generating bugs, which theoretically have seen many real-world bugs in its training data. This can potentially help BugFarm bugs be similar to real-world bugs compared to $\mu$BERT, which does not concern such similarity. We have identified all the potential threats to the validity in the design of the evaluation and have addressed them throughout the evaluation section when it was the most relevant. Our next step would be investigating the extent to which BugFarm bugs can enhance downstream applications of synthetic bugs in learning-based techniques, such as test case prioritization and bug localization.
6. Related Work
Our research is related to prior work on (1) real-world bug/vulnerability benchmarks, (2) mutation testing, and (3) learning-based bug/vulnerability generation.
Real-world bug/vulnerability benchmarks. Several attempts have been made to construct real-world bug datasets manually. Defects4J (Just et al., 2014a), BugSwarm (Tomassi et al., 2019), Bugs.jar (Saha et al., 2018), and RegMiner (Song et al., 2022) are the commonly used Java bug datasets that have mined GitHub to collect regression bugs from bug-fixing commits. BigVul (Fan et al., 2020) and CVEFixes (Bhandari et al., 2021) are real-world examples of vulnerabilities collected from bug-fixing reports in the CVE/NVD database (Booth et al., 2013). BugFarm complements the bugs and vulnerabilities in these datasets with hard-to-repair and hard-to-detect bugs. Also, the bugs in these datasets only represent human mistakes, which could be potentially different from the mistakes AI programming tools make.
Mutation testing. Mutation testing has been widely used in testing programs written in different languages (Deng et al., 2017; Schuler and Zeller, 2009; Just, 2014; Coles et al., 2016; Delamaro et al., 2001; Ma et al., 2005; Mirshokraie et al., 2013; Deng et al., 2015; Black, 2017; Okun et al., 2013), as well as testing program properties such as specifications (Martin and Xie, 2007), memory usage (Wu et al., 2017), and energy consumption (Jabbarvand and Malek, 2017). Syntactic mutation operators, where small syntactic changes in the code generate the artificial bugs, are not representative of real bugs (Just et al., 2014b). Constructing defect-based mutation operators alleviates the mentioned limitation but requires domain knowledge. If done manually, it also can be labor-intensive and time-consuming. BugFarm is an automated technique for generating bugs; hence, it is superior to non-automated mutation testing. The current implementation of BugFarm does not generate specific categories of bugs. However, by incorporating some domain knowledge as additional contexts to our prompts, one can expand the bug injection component of BugFarm to generate specific categories of bugs.
Learning-based bug/vulnerability generation. Learning-based bug generation techniques were first proposed to overcome the limitations of manual defect model construction in mutation testing (Brown et al., 2017). Such techniques leverage machine learning to learn the bug or vulnerability patterns from real-world bug fixes and generate mutants accordingly. DeepMutation (Tufano et al., 2019) is such a technique that relies on sequence-to-sequence neural machine translation for learning and generating bugs. SemSeed (Patra and Pradel, 2021) is a technique that extracts bug patterns from real-world bug fixes and injects them into other programs so that the bug in the new program is syntactically different but semantically similar. MutationMonkey (Beller et al., 2021) mines bug patterns from historical changes and transforms them into mutation operators semi-automatically.
VulGen (Nong et al., 2023) combines pattern mining and deep learning to generate realistic bugs. Specifically, it mines a large corpus of vulnerability-fixing commits to extract bug patterns, and trains a deep learning model on the same dataset to learn where to inject the bugs. LEAM (Tian et al., 2022) learns to mutate code from large examples of real-world bug-fixing commits. To generate syntactically correct bugs, LEAM represents code as a sequence of AST nodes and learns to apply grammar rules to select and modify the code. Finally, $\mu$BERT (Khanfir et al., 2023) produces buggy versions by replacing code tokens with the spacial <mask> token, and uses CodeBERT to predict the masked token. Both LEAM and $\mu$BERT incorporate beam search (Freitag and Al-Onaizan, 2017) to generate bugs that involve more than one statement.
BugFarm is superior to prior learning-based bug-generation techniques in several ways. First and foremost, BugFarm does not involve any training or fine-tuning effort to learn bug patterns and generate bugs. Consequently, it is independent of existing real-world bug datasets or a corpus of bug-fixing commits. Second, while the majority of prior work only generates one-line bugs, BugFarm can be configured to generate bugs that involve multiple statements. Third, BugFarm is the first technique that targets the generation of bugs that can challenge learning-based bug detectors and repair tools, or bugs that represent the AI programming tools’ mistakes, rather than human mistakes. Our empirical evaluation confirmed that these properties result in the generation of bugs that are hard-to-detect and hard-to-repair.
7. Concluding Remarks
Bug benchmarks are essential in software engineering to evaluate automated techniques concerning bugs. The advent of learning-based software engineering demands even more for automated bug-generation techniques, not only for evaluation but also for training or fine-tuning ML4SE models. In this paper, we presented BugFarm, a model-in-the-loop technique for the automated generation of hard-to-detect and hard-to-repair bugs. Our empirical evaluation shows the superiority of BugFarm compared to alternative bug generation approaches in (1) generating unique and high-quality bugs, (2) constructing bug benchmarks that are complex both in terms of detection and repair, and (3) injecting hard-to-detect bugs into arbitrary code. BugFarm does not rely on existing bug datasets and is model- and programming language-agnostic.
We believe our novel perspective on the bug generation problem offers several directions for future work. First of all, the focus of this paper is on attention analysis to identify the weak spots of transformer-based models. While state-of-the-art code-language models and LLMs are all based on transformers, we currently consider alternative approaches generalized to other neural architectures. Second, we plan to evaluate the quality of our bugs for improving other software engineering tasks, including bug localization, test oracles, and test-suite management. Based on the promising results in bug detection and repair, we are confident that our generated bugs are useful for better evaluation and challenging those tasks.